OpenStudy (anonymous):

Deon worked 16 hours last week at the grocery store and earned $124.00. If he continues to earn the same hourly pay, how many additional hours must he work to earn another $186.00?

OpenStudy (ahsome):

First find hourly rate: \[Hourly Rate=\frac{TotalRate}{HoursSpent}\] \[Hourly Rate=\frac{124}{16}\] \[Hourly Rate=$7.75\] Now take away the wanted amount from the amount he has \[186-124=$62\] Lastly, divide that number by hourly rate \[HoursNeeded=\frac{MoneyNeeded}{HourlyRate}\] \[HoursNeeded=\frac{62}{7.75}\] \[HoursNeeded=8\]

OpenStudy (kobeni-chan):

If you divide $124 by the 16 hours, you find out that he gets paid $7.75 per hour. If you divide 186 by 7.75, you'll get 24 as how many more hours he will need to work for an additional $186

OpenStudy (ahsome):

If this helped, please press the Best Response button. Thanks :)

OpenStudy (ahsome):

This is asuming you are using the before \(124\) dollars. If not, just divide 184 by 7.75